Can you help me solve this system by substitution?
y=3x
x+y=8

Answer by nabla(475)

Let's solve this with words at first:
y is the same as 3x.
Thus, x plus y must be x+3x. This is 4x. Thus, 4x must be 8. The value for x that makes this true is 2. Moreover, y must now be 3*2=6.
In symbols:
y=3x
x+y=8
-->
x+3x=8
4x=8
x=2
y=3(2)=6.